# An explicit local geometric Langlands for supercuspidal representations: the toral case

### v1

@@ Acknowledgement
The author thanks Konstantin Jakob and Daxin Xu for helpful discussions, and also thanks Zhiwei Yun for providing comments on a draft of the paper. The research of Lingfei Yi is supported by the grant No. JIH1414033Y of Fudan University.
@@ 3.1. Recollection on certain regular supercuspidal representations
We associate a Yu datum to $(S,\phi)$ as in [KalethaRegular, Proposition 3.7.8] with a slight variant. First, $S$ has a unique point $x$ in its Bruhat-Tits building, giving a point in the building of $G$. We can then define a sequence of root subsystems
@@
where $E_{r}^{\times}=1+u^{\lceil mr\rceil}\mathbb{C}[\![u]\!]$. By Lemma 3.6.1 of loc. cit., $\Phi_{r}$ is a Levi subsystem. Let $R=\{r_{(1)},r_{(2)},...,r_{(p)}\}$ be real numbers $r_{(1)}>r_{(2)}>\cdots>r_{(p)}>0$ where $\Phi_{r}$ takes a jump. Note $r_{(1)}=\nu$. Let $G_{i}^{\prime}\subset G(E)$ be the standard Levi subgroup containing $S_{E}$ with root system $\Phi_{r_{i}}$. Then $G_{i}=G(F)\cap G_{i}^{\prime}$ define twisted Levi subgroups of $G(F)$:
@@
Here two of $\phi_{i}$’s in loc. cit. are trivial because of our assumptions (24) and $\nu=r_{1}$. For $1\leq i\leq p$, $\phi_{i}$ is assumed to be trivial on $G_{i+1,sc}(F)$ where $G_{i+1,sc}$ is the simply-connected cover of the derived group of $G_{i+1}$, and $\phi_{i}$ has depth $r_{i}$. Moreover, $\phi_{i}$ needs to be $G_{i}$-generic in the sense of [Yu, §9], which we recall below.
@@
Denote by $\mathfrak{z}_{i,-r}^{*}$, $\mathfrak{g}_{i,-r}^{*}$, $\mathfrak{g}_{i-1,-r}^{*}$ the Moy-Prasad filtration on the dual spaces. For example, $\mathfrak{g}_{i,-r}^{*}$ is the subspace of $X\in\mathfrak{g}_{i}^{*}$ satisfying $\mathrm{ord}_{t}(X(\mathfrak{g}_{i,r+}))>0$. We say $\phi_{i}:G_{i}\rightarrow\mathbb{C}^{\times}$ is $G_{i}$-generic of depth $r_{(i)}$ if $\phi_{i}$ is trivial on $G_{i,r_{(i)}+}$, non-trivial on $G_{i,r_{(i)}}$, and its differential lands inside $\mathrm{d}\phi_{i}\in\mathfrak{z}_{i,-r_{(i)}}^{*}$ such that
(26) $\mathrm{d}\phi_{i}(H_{\alpha})=-r_{(i)},\qquad\forall\ \alpha\in\Phi(G_{i-1},S)\backslash\Phi(G_{i},S).$
@@
Let $(G_{i})_{x,r,r^{\prime}+}=\cup_{s>r^{\prime}}(G_{i})_{x,r,s}$. The quotient $V_{i}=(G_{i})_{x,r_{(i)},r_{(i)}/2}/(G_{i})_{x,r_{(i)},r_{(i)}/2+}$ is commutative. Extend $\phi_{i+1}$ to $G_{i}$ via projection with respect to root subspaces. The pairing $\langle a,b\rangle=\phi_{i+1}(aba^{-1}b^{-1})$ defines a symplectic form on $V_{i}$. Choose a Lagrangian subspace $\mathfrak{m}_{i}\subset V_{i}$. Denote by $(G_{i})_{x,r_{(i)},r_{(i)}/2,\mathfrak{m}_{i}}$ the preimage of $\mathfrak{m}_{i}$ in $(G_{i})_{x,r_{(i)},r_{(i)}/2}$  [^3] In the usual Yu’s construction, instead of making a Lagrangian, Heisenberg representation and Weil representation are used. We make this change so that the eventual representation is induced from a character. The same change is made in [BBMY, JKY]. Denote
@@
Consider character $\tilde{\phi}=\prod_{i=1}^{p}\phi_{i}$ on $J$. Here $\phi_{i}$ are extended to $J$ as follows. On $(G_{i})_{x,r_{(i)},r_{(i)}/2,\mathfrak{m}_{i}}\subset(G_{i})_{x,r_{(i)},r_{(i)}/2}$, for $j\leq i$ where $G_{j}\supset G_{i}$, $\phi_{j}|_{(G_{i})_{x,r_{(i)},r_{(i)}/2}}$ is restricted from $\phi_{j}$; for $j>i$ where $G_{j}\subset G_{i}$, the character $\phi_{j}$ is first extended to $G_{x,r_{(j)}/2+}$ via projection
$G_{x,r_{(j)}/2+}/G_{x,r_{(j)}+}\simeq\mathfrak{g}_{x,r_{(j)}/2+}/\mathfrak{g}_{x,r_{(j)}/2+}\rightarrow\mathfrak{g}_{x,r_{(j)}/2+}/\mathfrak{g}_{x,r_{(j)}/2+}\simeq G_{x,r_{(j)}/2+}/G_{x,r_{(j)}+},$
@@
$\mathfrak{g}(E)=\mathfrak{g}_{j}(E)\oplus\bigoplus_{\alpha\in\Phi(G,S)-\Phi(G_{j},S)}\mathfrak{g}_{\alpha}(E).$
Then restrict the above extension of $\phi_{j}$ to $(G_{i})_{x,r_{(i)},r_{(i)}/2,\mathfrak{m}_{i}}\subset(G_{i})_{x,r_{(i)}/2}\subset(G_{i})_{x,r_{(j)}/2+}\subset G_{x,r_{(j)}/2+}$.
@@ 3.2. Refined leading terms
Denote $\mathfrak{s}=\mathrm{Lie}(S)$, $\mathfrak{s}_{0}=\mathrm{Lie}(S_{0})=\mathrm{Lie}(S_{0+})$. The character $\phi$ induces a vector $\mathrm{d}\phi\in\mathfrak{s}^{*}$. As on the spectral side, denote
@@ 4.2.1. A normal subgroup
Consider $\tau_{i}=\mathfrak{g}(F)\cap(\bigoplus_{0\neq\alpha\in\Phi(G_{i},S)}\mathfrak{g}(E)_{\alpha})$, $11\leq i\leq p-$ 1. Define
(30) $\mathfrak{j}^{+}=\mathfrak{g}_{x,r_{(1)}+}+\sum_{i=1}^{p-1}(\tau_{i})_{r_{(i+1)}/2}+\sum_{i=2}^{p}(\mathfrak{a}_{i})_{r_{(i)}}.$
@@
(i): The linear form $\tilde{\phi}$ evaluated on $\mathfrak{j}$ is by first projecting to $\mathfrak{s}$ by quotienting out root subspaces, then applying $\phi$. The first step annihilates $\tau_{i}$. From the definition of $r_{(i)}$, the depth of $\phi$ on $\mathfrak{a}_{i}$ is $r_{(i)}$. The claim follows.
(ii): Observe $\mathfrak{j}=\mathfrak{j}^{+}+\mathfrak{s}_{0}$. It is easy to see that the adjoint action of $\mathfrak{g}_{x,r_{(1)}+}$ and $\mathfrak{s}_{0}$ preserve $\mathfrak{j}^{+}$. It remains to show that $[(\tau_{i})_{r_{(i+1)}/2},(\tau_{j})_{r_{(j+1)}/2}]\subset\mathfrak{j}^{+}$. Observe that $[(\tau_{i})_{r_{(i)}/2},(\tau_{j})_{r_{(j)}/2}]$ is spanned by vectors of the form $[\mathfrak{g}(E)_{\alpha,r_{(i+1)}/2},\mathfrak{g}(E)_{\beta,r_{(j+1)}/2}]$ for $\alpha\in\Phi(G_{i})$, $\beta\in\Phi(G_{j})$. Without loss of generality, we assume $i\leq j$, where $\tau_{i}\supset\tau_{j}$ and $r_{(i+1)}\geq r_{(j+1)}$.
If $\alpha+\beta\neq 0$ and is a root, then $\alpha+\beta\in\Phi(G_{i})$, so that
$[\mathfrak{g}(E)_{\alpha,r_{(i+1)}/2},\mathfrak{g}(E)_{\beta,r_{(j+1)}/2}]\subset\tau_{i}(E)_{r_{(i+1)}/2+r_{(j+1)}/2}\subset\tau_{i}(E)_{r_{(i+1)}/2}.$
If $\beta=-\alpha\in\Phi(G_{i})$, then $i=j$, and
$[\mathfrak{g}(E)_{\alpha,r_{(i+1)}/2},\mathfrak{g}(E)_{-\alpha,r_{(i+1)}/2}]\subset(\mathfrak{s}\cap\mathfrak{g}_{i}^{\mathrm{der}}(E))_{r_{(i+1)}}.$
Recall $\mathfrak{a}_{s}=\mathfrak{z}_{s}\cap\mathfrak{g}_{s-1}^{\mathrm{der}}$. Since $\mathfrak{s}\cap\mathfrak{g}_{i}^{\mathrm{der}}=\oplus_{s=i+1}^{p}\mathfrak{a}_{s}$, $(\mathfrak{s}\cap\mathfrak{g}_{i}^{\mathrm{der}}(E))_{r_{(i+1)}}\subset\oplus_{s=i+1}^{p}(\mathfrak{a}_{s})_{r_{(s)}}$. Summing up the above vectors, we conclude $[(\tau_{i})_{r_{(i+1)}/2},(\tau_{j})_{r_{(j+1)}/2}]\subset\mathfrak{j}^{+}(E)\cap\mathfrak{g}(F)=\mathfrak{j}^{+}$. ∎
@@ 6.1. Recollection on toral supercuspidal representations
$(A,B)\in\tau_{\frac{N}{2}}\times\tau_{\frac{N}{2}}\rightarrow\langle Y,[A,B]\rangle.$
Let $\mathfrak{m}\subset\tau_{\frac{N}{2}}$ be a Lagrangian subspace and $P(\frac{N}{2})_{\mathfrak{m}}$ the preimage of $\mathfrak{m}$ under $P(\frac{N}{2})/S(\frac{N}{2})P(\frac{N}{2}+1)\simeq\tau_{\frac{N}{2}}$. Let $J=S(1)P(\frac{N}{2})_{\mathfrak{m}}$ and let $\tilde{\phi}:J\rightarrow\mathbb{C}^{\times}$ be the composition of $\phi$ with projection to $S(1)$.
@@ 6.2. Segal-Sugawara operators
To write down $S_{i,j}$ more explicitly, we fix a topological basis of $\widehat{U}(\mathfrak{g})$ as follows. First we fix a topological basis of $\mathfrak{g}(\!(t)\!)$. Recall the (completed) grading given by (46). Each $\mathfrak{g}(\!(t)\!)_{i}\simeq\mathfrak{g}_{i}u^{i}=\mathfrak{t}_{Y,i}u^{i}\oplus\tau_{i}u^{i}$. Fix arbitrary basis $A_{i,k}$, $1\leq k\leq\dim\mathfrak{t}_{Y,i}$ of $\mathfrak{t}_{Y,i}$. For $\tau_{i}$, for each $\Theta$-orbit $O$ on $\Phi(\mathfrak{g},\mathfrak{t}_{Y})$, pick a representative $\alpha_{O}\in O$. Fix a basis $X_{\alpha_{O}}\in\mathfrak{g}_{\alpha_{O}}$. Then a basis of $\tau_{i}$ can be given by

### v2

@@ Acknowledgement
The author thanks Konstantin Jakob and Daxin Xu for helpful discussions, thanks Zhiwei Yun for providing comments on a draft of the paper, and thanks Will Sawin for pointing out a mistake in §3.1.2. The research of Lingfei Yi is supported by the grant No. JIH1414033Y of Fudan University.
@@ 3.1. Recollection on certain regular supercuspidal representations
We associate a Yu datum to $(S,\phi)$ as in [KalethaRegular, Proposition 3.7.8]. We slightly modify the notations to simplify them in our setting and to be more consistent with the notations on the spectral side. First, $S$ has a unique point $x$ in its Bruhat-Tits building, giving a point in the building of $G$. We can then define a sequence of root subsystems
@@
where $E_{r}^{\times}=1+u^{\lceil mr\rceil}\mathbb{C}[\![u]\!]$. By Lemma 3.6.1 of loc. cit., $\Phi_{r}$ is a Levi subsystem. Let $R=\{r_{(1)},r_{(2)},...,r_{(p)}\}$ be real numbers $r_{(1)}>r_{(2)}>\cdots>r_{(p)}>0$ where $\Phi_{r}$ takes a jump. Note $r_{(1)}=\nu$. Let $G_{i}^{\prime}\subset G(E)$ be the standard Levi subgroup containing $S_{E}$ with root system $\Phi_{r_{(i)}}$. Then $G_{i}=G(F)\cap G_{i}^{\prime}$ define twisted Levi subgroups of $G(F)$:
@@
Here two of $\phi_{i}$’s in loc. cit. are trivial because of our assumptions (24) and $\nu=r_{(1)}$. For $1\leq i\leq p$, $\phi_{i}$ is assumed to be trivial on $G_{i+1,sc}(F)$ where $G_{i+1,sc}$ is the simply-connected cover of the derived group of $G_{i+1}$, and $\phi_{i}$ has depth $r_{(i)}$. Moreover, $\phi_{i}$ needs to be $G_{i}$-generic in the sense of [Yu, §9], which we recall below.
@@
Denote by $\mathfrak{z}_{i,-r}^{*}$, $\mathfrak{g}_{i,-r}^{*}$, $\mathfrak{g}_{i-1,-r}^{*}$ the Moy-Prasad filtration on the dual spaces. For example, $\mathfrak{g}_{i,-r}^{*}$ is the subspace of $X\in\mathfrak{g}_{i}^{*}$ satisfying $\mathrm{ord}_{t}(X(\mathfrak{g}_{i,r+}))>0$. We say $\phi_{i}:G_{i}\rightarrow\mathbb{C}^{\times}$ is $G_{i-1}$-generic of depth $r_{(i)}$ if $\phi_{i}$ is trivial on $G_{i,r_{(i)}+}$, non-trivial on $G_{i,r_{(i)}}$, and its differential lands inside $\mathrm{d}\phi_{i}\in\mathfrak{z}_{i,-r_{(i)}}^{*}$ such that
(26) $\mathrm{ord}(\mathrm{d}\phi_{i}(H_{\alpha}))=-r_{(i)},\qquad\forall\ \alpha\in\Phi(G_{i-1},S)\backslash\Phi(G_{i},S).$
@@
Let $(G_{i})_{x,r,r^{\prime}+}=\cup_{s>r^{\prime}}(G_{i})_{x,r,s}$. The quotient $V_{i}=(G_{i})_{x,r_{(i+1)},r_{(i+1)}/2}/(G_{i})_{x,r_{(i+1)},r_{(i+1)}/2+}$ is commutative. Extend $\phi_{i+1}$ to $G_{i}$ via projection with respect to root subspaces. The pairing $\langle a,b\rangle=\phi_{i+1}(aba^{-1}b^{-1})$ defines a symplectic form on $V_{i}$. Choose a Lagrangian subspace $\mathfrak{m}_{i}\subset V_{i}$. Denote by $(G_{i})_{x,r_{(i+1)},r_{(i+1)}/2,\mathfrak{m}_{i}}$ the preimage of $\mathfrak{m}_{i}$ in $(G_{i})_{x,r_{(i+1)},r_{(i+1)}/2}$  [^3] In the usual Yu’s construction, instead of making a Lagrangian, Heisenberg representation and Weil representation are used. We make this change so that the eventual representation is induced from a character. The same change is made in [BBMY, JKY]. Denote
@@
Consider character $\tilde{\phi}=\prod_{j=1}^{p}\phi_{i}$ on $J$. Here $\phi_{i}$ are extended to $J$ as follows.
On $(G_{j})_{x,r_{(j+1)},r_{(j+1)}/2,\mathfrak{m}_{j}}\subset(G_{j})_{x,r_{(j+1)},r_{(j+1)}/2}$ where $i\leq j$, $G_{i}\supset G_{j}$, $\phi_{i}$ is restricted from $G_{i}$.
For $i>j$ where $G_{i}\subset G_{j}$, the character $\phi_{i}$ is first extended to $G_{x,r_{(i+1)}/2+}$ via projection
$G_{x,r_{(i+1)}/2+}/G_{x,r_{(i+1)}+}\simeq$ $\mathfrak{g}_{x,r_{(i+1)}/2+}/\mathfrak{g}_{x,r_{(i+1)}+}$
$\rightarrow$ $(\mathfrak{g}_{i})_{x,r_{(i+1)}/2+}/(\mathfrak{g}_{i})_{x,r_{(i+1)}+}\simeq(G_{i})_{x,r_{(i+1)}/2+}/(G_{i})_{x,r_{(i+1)}+},$
@@
$\mathfrak{g}(E)=\mathfrak{g}_{i}(E)\oplus\bigoplus_{\alpha\in\Phi(G,S)-\Phi(G_{i},S)}\mathfrak{g}_{\alpha}(E).$
Then restrict the above extension of $\phi_{i}$ to $(G_{j})_{x,r_{(j+1)},r_{(j+1)}/2,\mathfrak{m}_{j}}\subset(G_{j})_{x,r_{(j+1)}/2}\subset(G_{j})_{x,r_{(i+1)}/2+}\subset G_{x,r_{(i+1)}/2+}$.
@@ 3.2. Refined leading terms
Denote $\mathfrak{s}=\mathrm{Lie}(S)$, $\mathfrak{s}_{0}=\mathrm{Lie}(S_{0})=\mathrm{Lie}(S_{0+})=\mathfrak{s}_{0+}$. The character $\phi$ induces a vector $\mathrm{d}\phi\in\mathfrak{s}^{*}$. As on the spectral side, denote
@@ 4.2.1. A normal subgroup
Consider $\tau_{i}=\mathfrak{g}(F)\cap(\bigoplus_{0\neq\alpha\in\Phi(G_{i},S)}\mathfrak{g}(E)_{\alpha})$, $10\leq i\leq p-$ 1. Define
(30) $\mathfrak{j}^{\prime}=\mathfrak{g}_{x,r_{(1)}+}+\sum_{i=0}^{p-1}((\tau_{i})_{r_{(i+1)}/2}\cap(\mathfrak{g}_{i})_{x,r_{(i+1)},r_{(i+1)}/2,\mathfrak{m}_{i}})+\sum_{i=2}^{p}(\mathfrak{a}_{i})_{r_{(i)}+}.$
Let $\mathfrak{j}^{+}$ be the ideal of $\mathfrak{j}$ generated by subspace $\mathfrak{j}^{\prime}$. Denote the preimage of $\mathfrak{j}^{+}/\mathfrak{g}_{x,r_{(1)}}\subset\mathfrak{j}/\mathfrak{g}_{x,r_{(1)}}\simeq J/G_{x,r_{(1)}}$ in $J$ by $J^{+}$.
@@
The linear form $\tilde{\phi}$ evaluated on $\mathfrak{j}^{\prime}$ is by first projecting to $\mathfrak{s}_{0}$ by quotienting out root subspaces, then applying $\phi$. The first step annihilates $\tau_{i}$. From the definition of $r_{(i)}$, the depth of $\phi$ on $\mathfrak{a}_{i}$ is $r_{(i)}$. Since $\tilde{\phi}$ is the differential of a character, thus a Lie algebra homomorphism, it is then trivial on the ideal generated by $\mathfrak{j}^{\prime}$. ∎
Remark 13.
We expect that one can choose Lagrangians $\mathfrak{m}_{i}\subset V_{i}$ appropriately so that $\mathfrak{j}^{\prime}$ is itself an ideal of $\mathfrak{j}$. See (50) and Lemma 28 for the case of $p=1$.
@@ 6.1. Recollection on toral supercuspidal representations
Here although the notations $\mathfrak{g}_{i},\tau_{i}$ coincide with the Lie algebras of twisted Levi subalgebras in a general Yu datum, this should raise no confusion.
@@
(48) $(x,y)\in\tau_{\frac{N}{2}}\times\tau_{\frac{N}{2}}\rightarrow\langle Y,[x,y]\rangle.$
We choose a Lagrangian subspace $\mathfrak{m}\subset\tau_{\frac{N}{2}}$ as follows. Recall that the order $m$ automorphism $\Theta$ is regular elliptic, so that the subgroup of automorphisms generated by $\Theta$ acts freely on $\Phi(\mathfrak{g},\mathfrak{t}_{Y})$. For each $\Theta$-orbit $O\subset\Phi(\mathfrak{g},\mathfrak{t}_{Y})$, pick a representative $\alpha_{O}\in O$. Fix a basis $X_{\alpha_{O}}\in\mathfrak{g}_{\alpha_{O}}$. Then a basis of $\tau_{\frac{N}{2}}$ can be given by
(49) $X_{O,\frac{N}{2}}=X_{\alpha_{O}}+\zeta_{m}^{-\frac{N}{2}}\Theta(X_{\alpha_{O}})+\zeta_{m}^{-N}\Theta^{2}(X_{\alpha_{O}})+\cdots+\zeta_{m}^{-(m-1)\frac{N}{2}}\Theta^{m-1}(X_{\alpha_{O}}).$
Here $\Theta^{j}(X_{\alpha_{O}})\in\mathfrak{g}_{\Theta^{j}(\alpha_{O})}$. We claim that for any root $\alpha$ and $j$, $\Theta^{j}\alpha\neq-\alpha$. Otherwise $\Theta^{2j}\alpha=\alpha$, so that $m|2j$. Since $N$ is even, $(N,m)=1$, thus $m$ is odd, $m|j$, $\Theta^{j}\alpha=\alpha\neq-\alpha$.
Now we define a partition $\Theta\backslash\Phi(\mathfrak{g},\mathfrak{t}_{Y})=A\sqcup B$ of orbits as follows. Take any orbit $O$ and put it in $A$. For any $\alpha\in O$, from previous discussion we see $-\alpha\not\in O$, and we put the orbit of $-\alpha$ in B. Repeat this for the rest of orbits. The resulting partition satisfies $|A|=|B|=\frac{1}{2}\dim\tau_{\frac{N}{2}}$, and any $\alpha,\beta$ in the union of orbits in $A$ satisfy $\alpha\neq-\beta$. We let
(50) $\mathfrak{m}=\bigoplus_{O\in A}\mathbb{C}X_{O,\frac{N}{2}}.$
Thus $\dim\mathfrak{m}=\frac{1}{2}\dim\tau_{\frac{N}{2}}$. For any $O,O^{\prime}\in A$, since $\alpha\neq-\beta$ for any $\alpha\in O$ and $\beta\in O^{\prime}$, we have $[X_{O,\frac{N}{2}},X_{O^{\prime},\frac{N}{2}}]\in\tau\cap\mathfrak{g}_{N}=\tau_{N}$. Since the pairing of $\mathfrak{t}_{Y}$ with $\tau$ vanishes, $\mathfrak{m}$ is isotropic with respect to (48), thus a Lagrangian of $\tau_{\frac{N}{2}}$.
Let $P(\frac{N}{2})_{\mathfrak{m}}$ be the preimage of $\mathfrak{m}$ under $P(\frac{N}{2})/S(\frac{N}{2})P(\frac{N}{2}+1)\simeq\tau_{\frac{N}{2}}$. Let $J=S(1)P(\frac{N}{2})_{\mathfrak{m}}$ and let $\tilde{\phi}:J\rightarrow\mathbb{C}^{\times}$ be the composition of $\phi$ with projection to $S(1)$.
@@ 6.2. Segal-Sugawara operators
To write down $S_{i,j}$ more explicitly, we fix a topological basis of $\widehat{U}(\mathfrak{g})$ as follows. First we fix a topological basis of $\mathfrak{g}(\!(t)\!)$. Recall the (completed) grading given by (46). Each $\mathfrak{g}(\!(t)\!)_{i}\simeq\mathfrak{g}_{i}u^{i}=\mathfrak{t}_{Y,i}u^{i}\oplus\tau_{i}u^{i}$. Fix arbitrary basis $A_{i,k}$, $1\leq k\leq\dim\mathfrak{t}_{Y,i}$ of $\mathfrak{t}_{Y,i}$. For $\tau_{i}$, for each $\Theta$-orbit $O$ on $\Phi(\mathfrak{g},\mathfrak{t}_{Y})$, pick a representative $\alpha_{O}\in O$. Fix a basis $X_{\alpha_{O}}\in\mathfrak{g}_{\alpha_{O}}$. Similarly as for $i=\frac{N}{2}$, a basis of $\tau_{i}$ can be given by